OpenStudy (anonymous):

Solve the following equation; -4y+8=4(2y-2)-2(-16+8y)

OpenStudy (anonymous):

so \[-4y+8=4(2y-2)-2(-16+8y)\] use the distributive property to simplify:\[-4y +8 = 8y - 8 +32 - 16y\] add 4y to both sides:\[8 = 12y - 8 +32-16y\] Combine like terms:\[8 = -4y+24\] subtract 24 from both sides:\[-16 = -4y\] Divide by -4: \[4 = y\] SWITCHEROO!: \[y = 4\] ANSWER!!!!!!!: \[y = 4\] HOPE THIS HELPED! THX ;)